OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4> PLEASE HELP MEEE PLEASEE!!! :(

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (jdoe0001):

keep in mind that \(\huge {cos(\theta)=\frac{u \cdot v}{||u||\times ||v||}}\)

OpenStudy (jdoe0001):

once you get that value, just arcCosine it to get the angle between them

OpenStudy (anonymous):

I don't really know how to use even that??

OpenStudy (anonymous):

How do I know what u and v are?

OpenStudy (jdoe0001):

u = <6, -1>, v = <7, -4> <-----

OpenStudy (anonymous):

so I plug in both values? and solve twice?

OpenStudy (jdoe0001):

well, the numerator is the scalar value from the dot-product the denominator is the multiplication of their magnitudes

OpenStudy (jdoe0001):

<a, b> <c, d> dot product, as you should know is a*c+b*d

OpenStudy (jdoe0001):

and the magnitudes of <a, b> is just \(\bf \sqrt{a^2+b^2}\)

OpenStudy (anonymous):

ohh okay so how I set it up would be costheta= 6⋅7||6||×||7||?

OpenStudy (jdoe0001):

hmmm, let's first get the so-called "dot product" \(\bf <6, -1>\cdot <7, -4> \implies 6\times 7 + -1 \times -4\)

OpenStudy (jdoe0001):

now let's get the magnitudes \(\bf ||u|| = \sqrt{(6)^2+(-1)^2}\\ ||v|| = \sqrt{(7)^2+(-4)^2}\)

OpenStudy (jdoe0001):

so once you get those values, you use that to find the \(\bf cos(\theta)=\cfrac{u \cdot v}{||u||\times ||v||}\)

OpenStudy (anonymous):

let me grab my calcc

OpenStudy (anonymous):

so v=65

OpenStudy (anonymous):

and u=37

OpenStudy (jdoe0001):

well, square root of that, yes

OpenStudy (anonymous):

okay! after I have those two do I plug them in to the costheta formula?

OpenStudy (jdoe0001):

yes lemme check something

OpenStudy (anonymous):

okay :)

OpenStudy (jdoe0001):

well, I'm getting a number that doesn't work :/

OpenStudy (jdoe0001):

are the signs correct for => u = <6, -1>, v = <7, -4> ?

OpenStudy (anonymous):

yeah they are!! there are four answer choices too! 20.3° 10.2° 0.2° 30.3°

OpenStudy (anonymous):

if you got one of those it is probably correct

OpenStudy (jdoe0001):

ohh wait

OpenStudy (jdoe0001):

yes, I get a good number

OpenStudy (jdoe0001):

just had a wrong digit

OpenStudy (anonymous):

so just plug them into that?

OpenStudy (jdoe0001):

so the dot product is 46, thus $$\bf cos(\theta)=\cfrac{46}{\sqrt{37} \times \sqrt{65}}\\ $$

OpenStudy (jdoe0001):

then arcCosine them both sides $$ cos^{-1}(cos(\theta))=cos^{-1}\left(\cfrac{46}{\sqrt{37} \times \sqrt{65}}\right)\\ \theta = 20.28^o $$

OpenStudy (anonymous):

I GOT THATTTT!!!! :DDDD

OpenStudy (anonymous):

WOOOOOOO THANK YOUUUUUUU :) <3

OpenStudy (jdoe0001):

yw